### v1

# Vision-driven Autonomous Flight of UAV Along River Using Deep Reinforcement Learning with Dynamic Expert Guidance

Abstract
Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable river environment simulators and reward sparsity in such environments. To easily verify the navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity. Successful river following trajectories in the environment are manually collected and Behavior Clone (BC) is used to train an Imitation Learning (IL) agent to mimic expert behavior and generate expert guidance. Finally, a framework is proposed to train a Deep Reinforcement Learning (DRL) agent using BC expert guidance and improve the expert policy online by sampling good demonstrations produced by the DRL to increase convergence rate and policy performance. This framework is able to solve the along-river autonomous navigation task and outperform baseline RL and IL methods. The code and trainable environments are available [^1] https://github.com/lijianwen1997/Deep-Reinforcement-Learning-with-Dynamic-Expert-Guidance.
I INTRODUCTION
Autonomous flight of UAVs over inland waterways has been investigated in applications like water sampling [1], inspection of irrigation system [2], and cooperative drone-boat navigation tasks [3, 4, 5]. However, very few works addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways. Due to seasonal changes of river channels’ geometry and water level, manually planned GPS waypoints may not suit for accurate river following missions. Besides, when navigating in cluttered riverine environments, UAVs also need to avoid obstacles like trees and bridges, while being above the water.
To address the above challenges, a camera is used as the major perception source for the UAV to take advantage of information embedded in images and the recent development in image processing. To reduce the cost and safety issues for real-world experiments, we developed a series of riverine simulation environments using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input. The impact of environmental forces on UAV is not considered in the simulation environment and the UAV action in the 3D space is discretized to alleviate the navigation challenge of operating without a map along rivers.
Using the above simplification, we decouple the optimization of high-level local waypoint-based control policies or navigation planners for UAVs from mid-level velocity control, which is more suitable for the river following task due to its larger acceptable operation space compared to autonomous driving tasks for cars.
We pre-train an Imitation Learning (IL) agent using human demonstrations to provide basic guidance, then combine Reinforcement Learning (RL) algorithm with an online IL algorithm that accumulates and learns from decent agent experiences, to boost the learning speed and final agent performance. This will overcome the challenge of obtaining a performant controller using only RL algorithms since the only positive reward is given when the UAV traverses any newly visited discretized river segment. This training framework is examined to outperform Proximal Policy Optimization (PPO), Behavior Clone (BC), and PPO with static BC guidance.
In summary, the main contributions of this paper are:
  • The creation of the trainable photo-realistic riverine simulation environment, which bridges the sim2real gap for vision-driven autonomous navigation of UAVs in the river following task.
  • The development of Deep Reinforcement Learning with a dynamic expert guidance framework that interactively trains Reinforcement Learning and Imitation Learning agents to reduce the training time and increase the overall performance.
II RELATED WORK
Autonomous river following in a relatively simple curved river segment has been tested using a water segmentation mask from HSV threshed images [7, 8]. This image processing technique is preferable in terms of its implementation simplicity but has vulnerabilities like unverified generalization ability to rivers under different geometrical and light conditions, as well as the high safety hazards and experimental costs.
Game engines as simulation environments have brought more advantages for vision-driven Reinforcement Learning (RL) training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the sim2real gap in air [9, 10], land [11, 12] (indoor/outdoor), and even underwater applications [13].
To achieve a safer and more robust autonomous navigation controller in a diverse and adaptive riverine domain, researchers utilized Unreal Engine to train the agent for the river following task [14, 15]. Their proof of the superiority of the Variational Auto Encoder (VAE) [16] for image encoding, over Linear Regression method or end-to-end controller training, is adopted in our work. However, the reset criteria and the demonstrative trajectory collection approaches of their river simulation environment are manual-based, the agent action is limited to 2D planes, and the environment is not publicly available. To overcome these difficulties, our Unity-based river simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [17] environment, thanks to Unity ML-Agents Toolkit [6].
Reinforcement learning(RL) and Imitation Learning(IL) are two powerful techniques with significant applications in the field of robotics. Robots can learn optimal behavior through trial and error through RL. Robots can learn to avoid obstacles [18], plan paths [19], and perform tasks like autonomous driving [20] through RL even in complicated, dynamic situations. On the other hand, IL [21] leverages expert demonstrations to train robots. By monitoring and imitating the actions of experts, robots can acquire complex skills rapidly. This is valuable in scenarios where human expertise is readily available, as it allows robots to learn efficiently without extensive programming [22].
However, both of these approaches have some limitations. RL frequently requires a large number of interactions with the environment to learn optimal policies, which can be time-consuming and expensive. Recently, Researchers have combined RL and IL to accelerate the training of RL using IL’s guidance. For example, [23] proposed to use the imitative expert policy to guide the learning of the DRL agent by regularizing the divergence between the DRL agent’s policy and the expert policy. [24] proposed a framework that uses a loss function that combines behavior cloning and 1-step Q-learning losses to enable the transition from BC to RL without performance degradation. However, after the DRL agent has developed a policy that is similar to the expert policy if the DRL agent tries to update the policy to improve the performance, it will be constrained by the expert policy or the value penalty thus the improvement is hindered.
On the other hand, as they have predominantly learned to imitate the demonstrations, robots trained through IL may not be able to adapt well to unexpected circumstances. DAgger [25] improves on BC by training on a dataset that better resembles the observations the expert policy is likely to encounter. However, the IL agent is also constrained by the expert. Thus to overcome these limitations, we propose an interactive framework called DRL with dynamic expert guidance that trains an RL agent using expert guidance and improves the expert policy online by sampling demonstrations produced by the DRL expert so that IL and RL can be improved jointly.
(a) Comparison of image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
(b) Orthographic images of the training (left) and testing (right) river simulation environments.
Figure 1: Real and simulation river environments.
III SIMULATION
III-A Photo-realistic River Environments
The riverine simulation environment is built in Unity using River Auto Material 2019  [^2] https://assetstore.unity.com/packages/tools/terrain/r-a-m-2019-river-auto-material-2019-145937 and Mountain Environment  [^3] https://assetstore.unity.com/packages/3d/vegetation/mountain-environment-dynamic-nature-191851. The river spline tool is provided by the former package to support fast, flexible and scriptable river building with various realistic water textures and shaders. The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world river scene components and appearance, as shown in Figure 0(a).
Aside from the realistic river and terrain trees, the riverine simulation environment features several additional aspects that not only mimic the general objects in real-world rivers, creeks, and streams, but also make the training of river following agents more difficult: 1) river channels with varied widths and depths, 2) river tributaries that diverge from the main water stream, 3) bridges over the river that can obstruct the navigable path of UAVs, and 4) sharp and slow turns. (Figure 2)
Several environment design choices are made for better photo-realism and more robust training. Instead of continuous velocity control, discrete position control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc. Thus we will name the UAV agent as the camera agent in the contents hereafter. The circular shape of rivers in our simulated environments is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river. Besides, in the circular river environment, no matter where the camera agent is reset at the beginning of a training episode, the agent policy always has a chance to learn to traverse the whole river before ending this episode, making the collected trajectory lengths more uniform and independent of agent reset positions.
(a) Varying widths and depths
(b) Tributary
(c) Bridge
(d) Challenging U turns
Figure 2: Key features of the Unity river simulation environment.
TABLE I: Multi-discrete action values of the camera agent in the actual execution order, whose controlling logic is the same as remote controllers of common UAVs.
Action 0 1 2
Up-down Translation (m) 1 0 -1
Horizontal Rotation (deg) 10 0 -10
Forward-backward Translation (m) 1 0 -1
Leftward-rightward Translation (m) 0.5 0 -0.5
III-B Camera Agent
The simulation agent, created using Unity ML-Agents Toolkit [6], is a camera that looks forward with a fixed pitch angle (20 degrees in our experiments) with respect to the drone frame. Since no dynamics is included in this simulation, and all control actions are discrete, the drone camera itself is viewed as the agent. The camera agent is controlled by a four-dimensional multi-discrete action that maps from the command inputs of the common remote controller of UAVs: 1) translation up and down - throttle, 2) horizontal rotation counter-clockwise and clockwise - yaw rate, 3) translation forward and backward - pitch, 4) translation leftward and rightward - roll. Each action branch has three discrete values $\{0,1,2\}$, where $1$ means no action on that branch. The actions are applied in the drone coordinate frame (instead of the camera frame) with the right-hand rule, where the positive X axis points along the drone head and the positive Z axis points upward to the sky. The specific values of movement caused by actions are in Table I.
III-C Gym Environment
A river can be abstracted as a 3D Catmull-Rom spline [26] that is interpolated from several control points. The river surface is formed if we stretch this spline bi-directionally along some direction. This closed spline is lifted up to form a 3D bounding volume to limit the height of the camera agent, as well as its horizontal position. The tangents along the spline provide failure judgment when the camera agent’s yaw angle is far from either tangent angle of the nearest point on this spline with the projected agent position. This prohibits the turnaround motion of the agent since the built river is circular so that the agent can finish the whole-river navigation without any turnaround. Even if the agent does not trigger the above two ”done” criteria, lingering around or doing continuous meaningless movements with no reward improvement (over 50 steps in our experiments) will also trigger environment reset, and so does collision with bridges. If the environment ”done” is triggered during training, the camera agent will be reset at a random point inside the spline-based bounding volume, with a random yaw near either spline tangent of the nearest spline point. The above criteria for the river following task can be better visualized in Figure 5.
Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task. $h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments. The drone picture is from  [^5] https://www.jaycar.com.au/r-c-fpv-drone-with-1080p-camera/p/GT4118.
The river central spline can be simplified as a sequence of connected line segments with nearly the same length. To have a decent baseline RL algorithm to compare, and to intuitively sense the learning capabilities of different learning algorithms, shaped reward is given to the camera agent if the projection of its positions on the spline before and after an action covers any un-visited line segments. The line segment where the camera agent’s current position projects onto is denoted as visited. If the camera agent has visited all the river splines, it will get the highest reward, then the environment will be reset. For collision detection, a mesh collider is used for bridges, and a box collider with a side length of $0.5m$ is used for the camera agent. The overall reward function for the autonomous river following task is given below.
$r(s,a)=\begin{cases}-1,&\text{agent out of spline bounding volume}\\ -1,&\text{agent yaw diverts too much}\\ -1,&\text{agent collides with bridge}\\ -1,&\text{agent no progress over certain steps}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
where $r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case.
$r_{shaped}=10*n/N\$ (2)
All negative reward conditions denote the reset of the environment (Equation 1). Besides, no final positive reward is given when the camera agent has visited all river spline segments since the amplified shaped reward, Equation 2, already gives informative guidance for the river following task. Due to the same reason, no environment reset is applied for backup movement or staying still of the agent.
IV METHODOLOGY
Figure 4: System architecture. A human expert collects good trajectory before the training and the transitions are represented by blue arrows. $a^{H}_{t}$ denotes the human expert’s action. The transitions during training are represented by black arrows. $a^{E}_{t}$ denotes the IL expert’s action.
The proposed method is shown in Figure 4 and algorithm 1 to solve the river following task. It requires a demonstration dataset from a human expert, an IL agent to provide guidance, and an RL agent that explores the environment, updates the dataset and learns a control policy.
To train an IL agent to learn from demonstration, a dataset $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories is needed. Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...s_{t},a_{t}}$. The goal of the expert policy $\pi^{E}$ is to maximize the log-likelihood over $\mathcal{D}^{E}$.
[…]
To keep improving the expert policy’s performance during the DRL training, trajectories with episodic rewards larger than the reward threshold $T_{reward}$ will be appended to $\mathcal{D}^{E}$. $T_{reward}$ is $4$ in our experiments. Once the dataset is updated, the expert policy is retrained using the newest $2000$ steps from the dataset until it can mimic at least $95\%$ behaviors from the demonstration.
Algorithm 1 DRL with Dynamic Expert Guidance
Expert collects dataset $\mathcal{D}^{E}$
Train the IL policy $\pi^{E}$
Initialize policy network $\phi$, value network $\psi$
repeat
     Observe state $s_{t}$
     Sample actions $a_{t}\sim\pi^{\phi}$ and $a_{t}^{E}\sim\pi^{E}$
     Execute the action $a_{t}$ in the environment
     Observe the next state $s_{t+1}$ and reward $r_{t}$
     Store transition $(s_{t},a_{t},s_{t+1},r_{t})$ in the agent buffer $\mathcal{D}^{R}$
     if time to update DRL then
         Sample a mini-batch $\mathcal{B}$
         Calculate KL divergence $D_{KL}$
         Calculate loss $\mathcal{L}$
         Update value network $\psi$
         Update policy network $\phi$
     end if
     if time to evaluation DRL then
         Sample trajectories using $\pi^{\phi}$
         if $\tau$ has a reward above $T_{reward}$  then
              Append the trajectory $\tau$ in $\mathcal{D}^{E}$
              Update the IL policy $\pi^{E}$
         end if
     end if
until Convergence or maximum training steps
The target of RL is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
[…]
where $\tau$ is a trajectory and $p_{\pi}(\tau)$ is the distribution of the trajectory under policy $\pi$, and $T$ is the time horizon. Due to Proximal Policy Optimization(PPO)’s scalability and generalization ability, we use PPO [27] as a DRL algorithm to update the parameters of the policy network. In our framework, we use policy loss, value loss, and additional action loss to calculate the total loss and optimize the learning parameters.
The policy loss constrains the biggest possible improvement step on a policy using the data we currently have.
$\mathcal{L}_{policy}=\mathbb{E}\Bigl{[}\mathrm{min}{\Bigl{(}r_{t}(\theta)\hat{A_{t}},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A_{t}}\Bigr{)}}\Bigr{]}$ (5)
where $r_{t}(\theta)=\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{\theta}_{old}(a_{t}|s_{t})}$ denotes the probability ratio between the current and old policy. $\hat{A_{t}}$ is the advantage. The value function network $V_{\psi}$ is updated by the following loss function:
$\mathcal{L}_{value}=\mathbb{E}\Bigl{[}\Bigl{(}R_{t}-V_{\phi}(s_{t})\Bigr{)}^{2}\Bigr{]}$ (6)
where $R_{t}$ represents the returned value. To encourage the learned policy to be close to the policy learned from the demonstration, we can set a penalty term in the loss function
$\mathcal{L}_{action}=D_{KL}(\pi^{\theta},\pi^{E})=\sum\pi^{\theta}(a_{t}|s_{t})\log\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{E}(a_{t}|s_{t})}$ (7)
where $D_{KL}$ is the KL divergence that measures the divergence between two distributions.
Combining the above loss functions for the DRK with dynamic expert guidance becomes
$\mathcal{L}=c_{policy}\mathcal{L}_{policy}+c_{value}\mathcal{L}_{value}+c_{action}\mathcal{L}_{action}$ (8)
V SIMULATION VERIFICATION
One training and one testing river simulation environment 0(b) are created to train learning algorithms and test their generalization abilities. Specifically, these maps differ in the overall river length, variation of river widths and depths, number of (sharp) turns, bridges, tributaries, and sand islands, whereas the lighting conditions and terrain textures are unified to compare the performance of vision-driven controllers in learning geometrical patterns for the river following task.
Observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers. Variational Autoencoder [16] is obtained for each map by training 2000 randomly collected images from all maps up to the current training map’s difficulty level. This progressive update of the image encoder depends on the texture coherence across all maps, reduces the impact of the encoding module on RL training, and accounts for the downward compatibility of the trained agent. The encoded mean vector (without Gaussian noise) is used as the observation input to RL algorithms.
We ask an expert to execute the designed river following tasks in the simulator and collect actions and states as expert demonstration data. The UAV is randomly initialized in the valid activity space. The expert tries to correct the pose of the UAV and then controls the UAV to follow the river. Each trajectory is $50$ steps long. We collect $50$ trajectories in the medium environment as $\mathcal{D}^{E}$. The expert policy $\pi^{E}$ is trained by behavior clone using $\mathcal{D}^{E}$. After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.50$ in the training environment. The expert policy is used to generate reference actions during the training of RL agents.
Training is done on a desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU. Code related to RL and IL is based on StableBaseline3 [28] and Imitation Learning Baseline Implementations [29].
VI RESULTS AND ANALYSIS
VI-A Training Results
The proposed methods together with baseline methods (BC, PPO, and PPO+StaticBC) are used to train a camera agent in the river following environment. All the methods have been trained for $14000$ steps for $2$ different seeds.
Figure5 shows the training results of different methods in the training environment. The training curve is generated by averaging the training results from different seeds and represented by a solid line of the mean reward with ±$\pm 1$ standard error. The episodic reward is calculated by averaging $100$ past episodes and updated every $1024$ steps. The blue horizon line represents the average of $100$ episodic rewards of the expert policy, which is obtained from 50 episodes of evaluation in $2$ different seeds.
PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline. The PPO+DaynamicBC initially returns rewards similar to PPO+StaticBC baseline. Both methods can get a reward at or above the BC after $25000$ steps. However, after the DRL agent’s performance surpasses the BC, PPO+DaynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$. PPO+StaticBC also shows multiple decrease trends during training, mainly because of the constraint of the static expert policy. PPO+DaynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps. On average, PPO+DaynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than the BC, and $0.4$ higher than the PPO+StaticBC.
Figure 5: Training results of different approaches in the river following environment. The green horizontal line represents the performance of the pre-trained BC.
VI-B Testing Results
Figure 6: Episode reward distributions of four algorithms on training and testing environments. The horizontal bar represents the mean, the vertical bar represents 25% and 75% percentiles, and the white dot is the median. Statistics from 50 episodes for each algorithm.
Although both algorithms PPO+StaticBC and PPO+DynamicBC can achieve high episodic reward in the training map, the dynamic one still achieves higher mean and median episodic rewards in both maps and leads the static one in terms of the maximum episodic reward by a large margin in the testing map (Figure 6). This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to the PPO+StaticBC algorithm. PPO baseline and BC algorithms perform the worst in both training and testing maps, due to shortage of runtime expert guidance and limited generalization ability, respectively.
Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the training map, in two experiments. Star symbols signify the agent position every 50 steps. Trajectory color gets darker as the agent moves further away from the starting position.
To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the training map when starting at the same locations (Figure 7). The static algorithm agent can only navigate about $70$% of river length in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 0(b)) and bumping into the bridge, respectively. This is better visualized by the agent view near blue circles. The dynamic algorithm agent not only finishes whole-river navigation in both experiments but also with fewer steps, which justifies its robustness and efficiency. This can be visualized by the leading checkpoints (star in the figure) of the dynamic algorithm agent, meaning the agent moves a longer effective distance with fewer meaningless movements within the same 50 steps than the static agent. Besides, the dynamic algorithm handles better around river corners with smoother trajectory and larger safe margins to river banks. It also rejects tributary when seeing the main stream ahead by timely heading correction. Both cases can be viewed in figures near the red circles in Figure 7.
VII CONCLUSION
In this paper, we develop two riverine simulation environments to train a UAV agent to autonomously follow the river and avoid obstacles. A novel reinforcement learning method that utilizes the dynamic IL expert has been proposed. The performance of the proposed method has been compared with the baseline methods (PPO, BC, and PPO+StaticBC) in the riverine environments and achieves good qualitative and quantitative results. In the training environment, it achieves at least $0.4$ more rewards than all other methods on average during training and the highest best reward, and it is able to efficiently finish the river course without collision. It has good generalization ability as it also performs best in an unseen and difficult testing environment.
In the near future, tunable environmental light conditions like light directions and intensities can be enabled in the simulated riverine gym environment to increase policy robustness in the way of domain randomization and transfer the learned river following ability to a real UAV platform and test it in real-world riverine environment. On the other hand, we will systematically study and analyze the proposed reinforcement learning framework, keep optimizing the performance of the method, and apply it to solve more robotic challenges.